Please help me im failing and need help fast!
Solve the following systems:
1.)x/a+y=2b,
x/b-y=a-b

2.)ax-by=0,
x+y=a+b/ab
3.)x+b/a+y-b/b=a+b/b,
x-a/b-y-a/a=-a+b/a

Answer by pradhyumna.kumar@gmail.com(62)

Solve the following systems:
1.)x/a+y=2b,
x/b-y=a-b sol is : after adding the two equations
we get x = ab ,substitute the value of x in any of the equation we get y = b.
2.)ax-by=0,
x+y=a+b/ab sol : after adding the two equations we get a = 1/a and y = 1/b.
3.)x+b/a+y-b/b=a+b/b,
x-a/b-y-a/a=-a+b/a i am not able to get u'r question because typing fault i think so